Grade 11General Physics

The displacement x of a particle moving along x-axis at time t is given by x^2=2t^2+6t.The velocity at any time t is1). 4t+6/x2)2t-x/x3) 4t+64)2t+3/x

Anuradha
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
As we realize that ,
Velocity = dx/dt 
As it is given that 
x^2=2t^2+6t 
Velocity = d/dt(2t^2+6t) 
= 4t+6 m/s